# Towards Extreme Image Compression with Latent Feature Guidance and Diffusion Prior

### v2

@@ I Introduction
With the explosive growth of image data on the Internet, how to achieve efficient data transmission and storage has become increasingly important. Lossy image compression is a popular solution for saving storage and transmission bandwidth by exploiting the spatial and perceptual redundancy of images. Traditional compression standards, such as JPEG2000 [1], BPG [2], and VVC [3], are widely used in practice. However, these algorithms produce severe blocking artifacts at extremely low bitrates due to their block-based processing, see Fig. 1(b).
@@
3) We propose a conditional diffusion decoding module that fully exploits the powerful diffusion prior contained in the well-trained stable diffusion to facilitate extreme image compression and improve realistic reconstruction. In addition, we design a simple yet effective space alignment loss to enable end-to-end model training.
@@ II-A Lossy Image Compression
Lossy image compression plays a crucial role in image storage and transmission. Traditional compression standards, such as JPEG [24] and JPEG2000 [1], are widely used in practice. However, they tend to introduce block artifacts due to the lack of consideration of spatial correlation between image blocks. In recent years, learned image compression has made significant progress and achieved impressive rate-distortion performance [25, 26]. The main success of these methods is attributed to the development of various transform networks and entropy models. For instance, Liu et al. [27] introduce a non-local attention module to improve transform networks. In [28], He et al. employ invertible neural networks (INNs) to mitigate the information loss problem. Zhu et al. [29] construct nonlinear transforms using swin-transformers, achieving superior compression performance compared to CNNs-based transforms. In [18], Yang et al. innovatively use conditional diffusion models as decoders. Furthermore, several methods [6, 7, 30] enhance performance by improving entropy models. For example, Minnen et al. [30] combine hierarchical priors with autoregressive models to reduce spatial redundancy within latent features. In [31], He et al. assume the redundancy in spatial dimension and channel dimension is orthogonal and propose a multi-dimension entropy model. Qian et al. [32] utilize a transformer to enable entropy models to capture long-range dependencies. Guo et al. [26] explore capturing the dependencies along both the spatial and channel dimensions by using the causal global contextual prediction.
@@ II-B Extreme Image Compression
In some practical scenarios, such as underwater wireless communication, the bandwidth is too narrow to transmit the images or videos. To overcome this dilemma, extreme image compression towards low bitrates (e.g., below 0.1 bpp) is urgently needed. Several algorithms [10, 11, 33, 34, 35] leverage generative adversarial networks (GANs) for realistic reconstructions and bit savings. In [10], Agustsson et al. incorporate a multi-scale discriminator to synthesize details that cannot be stored at extremely low bitrates. Mentzer et al. [34] explore normalization layers, generator and discriminator architectures, training strategies, as well as perceptual losses, achieving visually pleasing reconstructions at low bitrates. However, these approaches suffer from the unstable training of GANs and inevitably introduce unpleasant visual artifacts.
In addition, some approaches use prior knowledge to achieve extreme image compression. Yue et al. [36] describe input images based on the down-sampled version and handcrafted features, and use these descriptions to reconstuct the images from a large-scale image database. Their method can achieve impressive compression performance when the large-scale image database contains images that are highly correlated with the input images. Gao et al. [37] leverage invertible networks, significantly reducing information loss during the compression process. Similarly, Wei et al. [38] employ invertible and generative priors to achieve extreme compression by rescaling images with extreme scaling factors (i.e., 16$\times$ and 32$\times$). In [39], Li et al. employ physical priors for extreme underwater image compression. Jiang et al. [40] utilize the text descriptions of images as prior to guide image compression. Inspired by the tremendous success of diffusion models in image generation, some methods [20, 21] use more powerful pre-trained text-to-image diffusion models as prior knowledge. In [20], Pan et al. encode images into short text embeddings and then generate high-quality images with pre-trained text-to-image diffusion models by feeding the text embeddings. Lei et al. [21] directly compress the short text prompts and binary contour sketches on the encoded side, and then use them as input to the pre-trained text-to-image diffusion model for reconstruction on the decoded side. However, these diffusion-based methods treat pre-trained text-to-image diffusion models as independent components, which limits their ability to fully exploit the generative capabilities of pre-trained diffusion models.
@@ III-A3 Network Details
Figure 3: The architecture of the proposed LFGCM. $Conv\ k3s1$ denotes convolution with $3\times 3$ filters and stride 1. $Tconv\ k3s1$ denotes transposed convolution with $3\times 3$ filters and stride 1. $RB$ denotes residual block [46]. $RBneck$ denotes residual bottleneck block [46]. $LReLU$ denotes the LeakyReLU function. $AE$ and $AD$ denote arithmetic encoder and decoder, respectively. $C_{m}$ denotes context model. The black and red arrows denote main and guidance flow, respectively.
@@ III-B2 Conditional Diffusion Decoding Module
The CDDM is designed to leverage the powerful generative capability of fixed stable diffusion to reconstruct image $x$ with realistic details at extremely low bitrates. Inspired by ControlNet [14], we introduce a control module to inject content information contained in $z_{c}$ into the denoising process. This control module has the same encoder and middle block architecture as the noise estimator $\epsilon_{\theta}$. Notably, we reduce the channel number of the control module to 20% of the original, which results in a slight performance decrease but significantly enhances inference speed (see Section V-B). In addition, we increase the channel number of the first convolution layer to 6 to accommodate the concatenated input of the content variable $z_{c}$ and the latent noise $z_{t}$. Through the control module, we obtain a series of conditional features that contain content information and align with the internal knowledge of stable diffusion. These conditional features are then added to the encoder and decoder of the noise estimator $\epsilon_{\theta}$ using 1$\times$1 convolutions. Leveraging the powerful generative capability encapsulated in pre-trained stable diffusion, we can obtain a high perceptual quality reconstruction $\hat{x}$ even at extremely low bitrates.
@@ IV Experiments
(a) Original (b) ELIC (c) HiFiC (d) MS-ILLM (e) Text+Sketch (f) Wei et al. (g) DiffEIC (Ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Figure 6: Visual comparisons of the proposed DiffEIC framework with the MSE-optimized ELIC [31], the GANs-based HiFiC [34] and MS-ILLM [35], the diffusion-based Text+Sketch [21], and the method by Wei et al. [38] on the CLIC2020 [49] dataset. For each method, the bpp and DISTS values are shown beneath images. Compared to other methods, our method produces more realistic and faithful reconstructions with lower bpp.
@@ IV-B Comparisons With State-of-the-art Methods
We compare our DiffEIC with state-of-the-art learned image compression methods, including ELIC [31], HiFiC [34], Text+Sketch [21], Wei et al. [38], and MS-ILLM [35]. In addition, we compare with traditional image compression methods BPG [2] and VVC [3]. For BPG software, we optimize image quality and compression efficiency with the following settings: “YUV444” subsampling mode, “x265” HEVC implementation, “8-bit” depth, and “YCbCr” color space. For VVC, we employ the reference software VTM-23.0222https://vcgit.hhi.fraunhofer.de/jvet/VVCSoftware_VTM/-/tree/master with intra configuration.
@@ IV-B1 Quantitative Comparisons
The rate-distortion performance comparison is shown in Fig. 5. Since Text+Sketch [21] has ignored the pixel-level fidelity of the reconstruction results, we do not report its rate-distortion performance. Compared to Wei et al. [38], the proposed DiffEIC achieves better PSNR and MS-SSIM values. However, we find that DiffEIC is worse than other comparative methods. The reason behind this is that the proposed DiffEIC uses stable diffusion prior for realistic detail reconstruction at extremely low bitrates, which does not ensure pixel-level accuracy. To further demonstrate this, we report the PSNR and MS-SSIM values of the stable diffusion autoencoder (see the black horizontal line in Fig. 5), which can be treated as the upper bound of the performance of DiffEIC. Although it sacrifices some fidelity, the proposed DiffEIC is able to capture realism at extremely low bitrates.
@@ IV-B3 Complexity Comparisons
We further compare the proposed DiffEIC with state-of-the-art image compression methods in terms of complexity. For PerCo [19], we directly show the results reported in their paper, since the source codes are not available. Table I summarizes the average encoding/decoding time in seconds with its standard deviation on the Kodak dataset. On the one hand, it is worth noting that the diffusion-based methods have higher encoding and decoding complexity than the VAE-based and GAN-based methods. On the other hand, the proposed DiffEIC encoder is significantly faster than Text+Sketch [21]. Compared to PerCo [19], the proposed DiffEIC is able to achieve comparable encoding speed and faster decoding speed with the same number of denoising steps.
@@ V Analysis and Discussions
Methods Hyper-Parameter BD-Rate (%) Speed (in sec.)
CN (%) Denoising Steps DISTS LPIPS Encoding Speed Decoding Speed
DiffEIC (CN) 100 50 -5.36 -2.57 0.128 $\pm$ 0.005 6.012 $\pm$ 0.012
50 50 -2.99 -2.28 5.068 $\pm$ 0.020
DiffEIC (W/o LFG) 20 50 23.88 13.19 0.062 $\pm$ 0.009 4.574 $\pm$ 0.006
DiffEIC (Ours) 20 50 0 0 0.128 $\pm$ 0.005 4.574 $\pm$ 0.006
DiffEIC (DS) 20 20 22.51 6.20 0.128 $\pm$ 0.005 1.964 $\pm$ 0.009
20 10 37.83 13.26 1.089 $\pm$ 0.009
20 5 49.93 21.68 0.646 $\pm$ 0.005
20 0 59.50 35.77 0.212 $\pm$ 0.006
@@ V-A Ablation of Latent Feature Guidance
Figure 9: Quantitative comparisons of different denoising steps on the CLIC2020 [49] dataset. 0 step denotes directly using the decode $\mathcal{D}$ to decode content variables.
@@ V-B Effect of the Channel Number in Control Module
We further analyze how the number of channels of the control module affects the performance and complexity of the proposed DiffEIC. In our default setting, we reduce the number of channels to 20% of the original. We also increase the number of channels by setting the percentage to 50% and 100%. As shown in Table II, using more channels is able to bring a slight improvement in performance, where the lower DISTS and LPIPS values are achieved. However, it inevitably leads to the decoding complexity up. For example, the decoding speed of the proposed DiffEIC with 100% channels decrease by about 31% compared to the default setting. To achieve a tradeoff between performance and inference speed, we choose the 20% in the proposed DiffEIC.

### v3

@@ I Introduction
Extreme image compression, which aims to compress images at bitrates below 0.1 bits per pixel (bpp), is critical in very bandwidth-constrained scenarios, such as satellite communications. Traditional compression standards, such as JPEG2000 [1], BPG [2], and VVC [3], are widely used in practice. However, these algorithms produce severe blocking artifacts at extremely low bitrates due to their block-based processing, see Fig. 1(b).
@@
3) We propose a conditional diffusion decoding module that fully exploits the powerful diffusion prior contained in the well-trained stable diffusion to facilitate extreme image compression and improve realistic reconstruction.
@@ II-A Lossy Image Compression
Lossy image compression plays a crucial role in image storage and transmission. Traditional compression standards, such as BPG [2], HEVC [24] and VVC [3], are widely used in practice. However, they tend to introduce block artifacts due to the lack of consideration of spatial correlation between image blocks. In recent years, learned image compression has made significant progress and achieved impressive rate-distortion performance [25, 26]. The main success of these methods is attributed to the development of various transform networks and entropy models. For instance, Liu et al. [27] introduce a non-local attention module to improve transform networks. In [28], He et al. employ invertible neural networks (INNs) to mitigate the information loss problem. Zhu et al. [29] construct nonlinear transforms using swin-transformers, achieving superior compression performance compared to CNNs-based transforms. In [18], Yang et al. innovatively use conditional diffusion models as decoders. Furthermore, several methods [6, 7, 30] enhance performance by improving entropy models. For example, Minnen et al. [30] combine hierarchical priors with autoregressive models to reduce spatial redundancy within latent features. In [31], He et al. assume the redundancy in spatial dimension and channel dimension is orthogonal and propose a multi-dimension entropy model. Qian et al. [32] utilize a transformer to enable entropy models to capture long-range dependencies. Guo et al. [26] explore capturing the dependencies along both the spatial and channel dimensions by using the causal global contextual prediction.
@@ II-B Extreme Image Compression
In some practical scenarios, such as satellite communications, the bandwidth is too narrow to transmit the images or videos. To overcome this dilemma, extreme image compression towards low bitrates (e.g., below 0.1 bpp) is urgently needed. Several algorithms [10, 11, 33, 34, 35] leverage generative adversarial networks (GANs) for realistic reconstructions and bit savings. In [10], Agustsson et al. incorporate a multi-scale discriminator to synthesize details that cannot be stored at extremely low bitrates. Mentzer et al. [34] explore normalization layers, generator and discriminator architectures, training strategies, as well as perceptual losses, achieving visually pleasing reconstructions at low bitrates. However, these approaches suffer from the unstable training of GANs and inevitably introduce unpleasant visual artifacts.
In addition, some approaches use prior knowledge to achieve extreme image compression. Yue et al. [36] describe input images based on the down-sampled version and handcrafted features, and use these descriptions to reconstuct the images from a large-scale image database. Their method can achieve impressive compression performance when the large-scale image database contains images that are highly correlated with the input images. Benefiting from the bijective and information-lossless property of invertible neural networks (INNs), Gao et al. leverage INNs to mitigate the significant information loss in extreme image compression [37]. Wei et al. employ invertible and generative priors to achieve extreme compression by rescaling images with extreme scaling factors (i.e., 16$\times$ and 32$\times$) [38]. In [39], Li et al. employ physical priors (i.e., attenuation coefficient and ambient light) and the semantic prior for extreme underwater image compression, which may not generalize well to natural images with different scenarios. Jiang et al. utilize text descriptions as prior to guide image compression for better compression performance [40].
Inspired by the tremendous success of diffusion models in image generation, some methods [20, 21] use more powerful pre-trained text-to-image diffusion models as prior knowledge. In [20], Pan et al. encode images into short text embeddings and then generate high-quality images with pre-trained text-to-image diffusion models by feeding the text embeddings. Lei et al. [21] directly compress the short text prompts and binary contour sketches on the encoded side, and then use them as input to the pre-trained text-to-image diffusion model for reconstruction on the decoded side. However, these diffusion-based methods treat pre-trained text-to-image diffusion models as independent components, which limits their ability to fully exploit the generative capabilities of pre-trained diffusion models.
@@ III-A3 Network Details
Figure 3: The architecture of the proposed LFGCM. $Conv\ k3s1$ denotes convolution with $3\times 3$ filters and stride 1. $Tconv\ k3s1$ denotes transposed convolution with $3\times 3$ filters and stride 1. $RB$ denotes residual block [46]. $RBneck$ denotes residual bottleneck block [46]. $LReLU$ denotes the LeakyReLU function. $AE$ and $AD$ denote arithmetic encoder and decoder, respectively. $C_{m}$ denotes context model. $Fusion$ denotes the fusion method. The black and red arrows denote main and guidance flow, respectively.
@@ III-B2 Conditional Diffusion Decoding Module
The CDDM is designed to leverage the powerful generative capability of fixed stable diffusion to reconstruct image $x$ with realistic details at extremely low bitrates. Inspired by ControlNet [14], we introduce a control module to inject content information contained in $z_{c}$ into the denoising process. This control module has the same encoder and middle block architecture as the noise estimator $\epsilon_{\theta}$. Notably, we reduce the channel number of the control module to 20% of the original, which results in a slight performance decrease but significantly enhances inference speed (see Section V-B). In addition, we increase the channel number of the first convolution layer to 8 to accommodate the concatenated input of the content variable $z_{c}$ and the latent noise $z_{t}$. Through the control module, we obtain a series of conditional features that contain content information and align with the internal knowledge of stable diffusion. These conditional features are then added to the encoder and decoder of the noise estimator $\epsilon_{\theta}$ using 1$\times$1 convolutions. Leveraging the powerful generative capability encapsulated in pre-trained stable diffusion, we can obtain a high perceptual quality reconstruction $\hat{x}$ even at extremely low bitrates.
[…]
@@ IV Experiments
(a) Original (b) ELIC (c) HiFiC (d) MS-ILLM (e) Text+Sketch (f) PerCo (g) Wei et al. (h) DiffEIC (Ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 6: Visual comparisons of the proposed DiffEIC framework with the MSE-optimized ELIC [31], the GANs-based HiFiC [34] and MS-ILLM [35], the diffusion-based Text+Sketch [21] and PerCo [19], and the method by Wei et al. [38] on the CLIC2020 [49] dataset. For each method, the bpp and DISTS values are shown beneath images. Compared to other methods, our method produces more realistic and faithful reconstructions with lower bpp.
@@ IV-B Comparisons With State-of-the-art Methods
We compare our DiffEIC with state-of-the-art learned image compression methods, including ELIC [31], HiFiC [34], Text+Sketch [21], PerCo [19], Wei et al. [38], and MS-ILLM [35]. For PerCo [19], we use PerCo(SD) [^2] https://github.com/Nikolai10/PerCo/tree/master as a substitute, since the official source codes and models are not available. In addition, we compare with traditional image compression methods BPG [2] and VVC [3]. For BPG software, we optimize image quality and compression efficiency with the following settings: “YUV444” subsampling mode, “x265” HEVC implementation, “8-bit” depth, and “YCbCr” color space. For VVC, we employ the reference software VTM-23.0333https://vcgit.hhi.fraunhofer.de/jvet/VVCSoftware_VTM/-/tree/master with intra configuration.
@@ IV-B1 Quantitative Comparisons
The rate-distortion performance comparison is shown in Fig. 5. Since Text+Sketch [21] has ignored the pixel-level fidelity of the reconstruction results, we do not report its rate-distortion performance. Compared to Wei et al. [38] and PerCo [19], the proposed DiffEIC achieves better PSNR and MS-SSIM values. However, we find that DiffEIC is worse than other comparative methods. The reason behind this is that the proposed DiffEIC uses stable diffusion prior for realistic detail reconstruction at extremely low bitrates, which does not ensure pixel-level accuracy. To further demonstrate this, we report the PSNR and MS-SSIM values of the stable diffusion autoencoder (see the black horizontal line in Fig. 5), which can be treated as the upper bound of the performance of DiffEIC. Although it sacrifices some fidelity, the proposed DiffEIC is able to capture realism at extremely low bitrates.
@@ IV-B3 Complexity Comparisons
We further compare the proposed DiffEIC with state-of-the-art image compression methods in terms of complexity. For PerCo [19], we directly show the results reported in their paper, since the official source codes are not available. Table I summarizes the average encoding/decoding time in seconds with its standard deviation on the Kodak dataset. On the one hand, it is worth noting that the diffusion-based methods have higher encoding and decoding complexity than the VAE-based and GAN-based methods. On the other hand, the proposed DiffEIC encoder is significantly faster than Text+Sketch [21]. Compared to PerCo [19], the proposed DiffEIC is able to achieve comparable encoding speed and faster decoding speed with the same number of denoising steps.
@@ V-A Ablation of Latent Feature Guidance
Methods Hyper-Parameter BD-Rate (%) Speed (in sec.)
CN (%) Denoising Steps DISTS LPIPS Encoding Speed Decoding Speed
DiffEIC (CN) 100 50 -5.36 -2.57 0.128 $\pm$ 0.005 6.012 $\pm$ 0.012
50 50 -2.99 -2.28 5.068 $\pm$ 0.020
100 37 4.11 0.80 4.578 $\pm$ 0.013
50 43 0.96 0.19 4.474 $\pm$ 0.019
DiffEIC (W/o LFG) 20 50 23.88 13.19 0.062 $\pm$ 0.009 4.574 $\pm$ 0.006
DiffEIC (Ours) 20 50 0 0 0.128 $\pm$ 0.005 4.574 $\pm$ 0.006
DiffEIC (DS) 20 20 22.51 6.20 0.128 $\pm$ 0.005 1.964 $\pm$ 0.009
20 10 37.83 13.26 1.089 $\pm$ 0.009
20 5 49.93 21.68 0.646 $\pm$ 0.005
20 0 59.50 35.77 0.212 $\pm$ 0.006
@@
We further explore different fusion methods in the proposed LFGCM, including the addition, concatenation, and cross-attention. As shown in Table III, the fusion method using addition causes a severe degradation in compression performance. In addition, using the cross-attention mechanism achieves better compression performance than our method, while it increases the computational complexity by 31% in terms of MACs. Considering the trade-off between performance and computational complexity, we choose the concatenation method for feature fusion in the proposed DiffEIC.
TABLE III: Ablation studies of the fusion method in LFGCM. BD-Rate is evaluated on CLIC2020 dataset and the MACs of different fusion methods are calculated based on 768$\times$512 image patch.
Methods BD-Rate (%) MACs (G)
DISTS LPIPS NIQE MS-SSIM
Addition 6.66 5.24 20.46 2.43 1.13 (-17%)
Cross-Attention -0.45 -0.67 -2.08 -1.58 1.79 (+31%)
Concatenation(Ours) 0 0 0 0 1.36 (+ 0%)
Figure 9: Quantitative comparisons of different denoising steps on the CLIC2020 [49] dataset. 0 step denotes directly using the decoder $\mathcal{D}$ to decode content variables.
@@ V-B Effect of the Channel Number in Control Module
We further analyze how the number of channels of the control module affects the performance and complexity of the proposed DiffEIC. In our default setting, we reduce the number of channels to 20% of the original. We also increase the number of channels by setting the percentage to 50% and 100%. As shown in Table II, using more channels is able to bring a slight improvement in performance, where the lower DISTS and LPIPS values are achieved. However, it inevitably leads to the decoding complexity up. For example, the decoding speed of the proposed DiffEIC with 100% channels decrease by about 31% compared to the default setting. When the decoding time is comparable, using 20% channels results in slightly better performance than using higher percentages of channels. This indicates that reducing the number of channels is more effective in balancing performance and inference speed than simply reducing the number of denoising steps. To achieve a tradeoff between performance and inference speed, we choose the 20% in the proposed DiffEIC.